ROTATIONAL ANALYSIS OF BANDS IN THE HIGH-RESOLUTION INFRARED SPECTRA OF trans, trans- AND cis, cis-1,4-DIFLUOROBUTADIENE-2-d_{1}

\underline{NORMAN C. CRAIG}, DEACON J. NEMCHICK, CLAY C. EASTERDAY, ETHAN C. GLOR, and DREW F. K. WILLIAMSON, Department of Chemistry and Biochemistry, Oberlin College, Oberlin, OH 44074; THOMAS A. BLAKE and ROBERT L. SAMS, Environmental Molecular Sciences Laboratory, Pacific Northwest National Laboratory, Richland, WA 99352.

Ground state rotational constants for a series of isotopomers are being sought for use in determining the semi-experimental equilibrium structures of the isomers of 1,4-difluorobutadiene. Because fluorine substitution has a large influence on CC bond lengths in C_{3} and C_{4} rings, we asked how fluorine substitution affects butadiene. trans,trans- and cis,cis-1,4-Difluorobutadiene-2-d_{1} have been synthesized, and high-resolution (0.0013 cm^{−1}) infrared spectra have been recorded for these nonpolar species. Analysis of the rotational structure in several bands is reported. For the trans,trans isomer, the C-type band at 709.0 cm^{−1} for ν_{21}(a′′) has been fully analyzed, and the C-type band at 914.3 cm^{−1} for ν_{18}(a′′) has been partially analyzed. Interfering with the analysis of the second band is overlap of its R branch with the P branch of the A/B-type band for ν_{13}(a′) at 933 cm^{−1}. For the cis,cis isomer, as much as possible of the C-type band (K′_{a} = 10 to 34) for ν_{20}(a′′) at 775.4 cm^{−1} has been analyzed. An A-type band for ν_{13}(a′) at 865.8 cm^{−1} has also been analyzed into the band center. Small inertial defects confirm that these molecules are planar. Ground state rotational constants are reported for both isomers in comparison with those for the normal species.^{a}

aN. C. Craig, M. C. Moore, C. F. Neese, D. C. Oertel, L. Pedraza, and T. Masiello, J. Mol. Spectrosc. 254, 39-46 (2009).